OpenStudy (anonymous):

What's the parabolic standard form of x^2 -12x - 48y -372 = 0

OpenStudy (campbell_st):

the standard form of a parabola is \[y = ax^2 + bx + c\] so rewrite you equation start by adding 48y to both sides of the equation. then you well need to divide each term by 48 to get a single y and subsequently, your equation in the form above.

OpenStudy (anonymous):

try making perfect squares with x as the term x^2 is present

OpenStudy (anonymous):

x^2 -12x - 48y -372 = 0 or hint x^2 -12x +36 = 48y +372 +36

OpenStudy (campbell_st):

wow I thought (x - h)^2 = 4a(y -k) was a vertex form

OpenStudy (campbell_st):

I suppose it depends on there you learn your maths

OpenStudy (ybarrap):

nope, you're right.

OpenStudy (amistre64):

parabola is a geometric term, and as such should prolly ustilize the geometric standard definition :)

OpenStudy (jdoe0001):

hmm, "(x - h)^2 = 4a(y -k)" standard focus form "y = (x-h)^2 + k" standard vertex form ax^2+b+c standard form

OpenStudy (ybarrap):

$$ \tt From~x^2-12x-48y-372=0,~we~get~that~y=\frac{x^2}{48}-\frac{x}{4}-\frac{31}{4}\\ Assuming~Standard~form~is~(x-h)^2=4p(y-k) (^\star)\\ Where~(h,k)~is~vertex~and~p~is~focus.\\ We~have, with~y=ax^2+bx+c~,k=\frac{-b}{2a}~,h=\frac{4ac-b^2}{4a}~and~p=\frac{1}{4a}.\\ Substitute~h,~k~and~p~into~(^\star)~and~you're~done.\\ I ~get~ k=6,~h = -8.5~ and ~p = 12. $$